# GeoBotsVR: A Robotics Learning Game for Beginners with Hands-on Learning Simulation

### 20 Feb 2024 v3

@@ 1. Introduction
In terms of mode of learning, researchers suggest that virtual reality (VR) enhances learning by increasing engagement and immersion (Hussein and Nätterdal, 2015; Dede, 2009; Mubarrat et al., 2020; Srinivasan et al., 2021; Chowdhury et al., 2022b, a). VR’s spatial navigation reduces the cognitive load in programming learning, surpassing traditional text-based methods (Lee and Wong, 2014). Additionally, VR provides a sense of self-presence (Ratan, 2013), fostering an embodied-cognitive learning experience (Melcer, 2018; Shin, 2017) that encourages more intuitive interaction with the content (Steed et al., 2016), potentially improving overall learning outcomes (Cheon and Grant, 2012). Therefore, I propose GeoBotsVR, an easily accessible and easy-to-play VR puzzle-solving/racing game designed to foster interest and motivation in robotics, electronics, and embedded systems of players through an enjoyable experience. In GeoBotsVR, the main gameplay focuses on players solving procedurally-generated puzzles using a customizable robot in different levels, while learning about robotics components, electronics, and programming along the way.
@@ 2.1. Gamified Approaches Towards Programming Learning
2.2. Gamified Approaches Towards Electronics Learning
@@ 3.2. The Workshop: Robot Component Editor
3.3. The Arcade: Main Gameplay

### 27 Feb 2024 v4

@@ 1. Introduction
In terms of mode of learning, researchers suggest that virtual reality (VR) enhances learning by increasing engagement and immersion (Hussein and Nätterdal, 2015; Dede, 2009; Mubarrat et al., 2020; Srinivasan et al., 2021; Mubarrat et al., 2022b, a). VR’s spatial navigation reduces the cognitive load in programming learning, surpassing traditional text-based methods (Lee and Wong, 2014). Additionally, VR provides a sense of self-presence (Ratan, 2013), fostering an embodied-cognitive learning experience (Melcer, 2018; Shin, 2017) that encourages more intuitive interaction with the content (Steed et al., 2016), potentially improving overall learning outcomes (Cheon and Grant, 2012). Therefore, I propose GeoBotsVR [^1] Preview video: https://youtu.be/Q0t3EfVHfrc [^2] Demo video: https://youtu.be/CBdWhMo7stQ, an easily accessible and easy-to-play VR puzzle-solving/racing game designed to foster interest and motivation in robotics, electronics, and embedded systems of players through an enjoyable experience. In GeoBotsVR, the main gameplay focuses on players solving procedurally-generated puzzles using a customizable robot in different levels, while learning about robotics components, electronics, and programming along the way.
@@ 2.1. Gamified Approaches Towards Programming Learning
$2.$ Gamified Approaches Towards Electronics Learning
@@ 3.2. The Workshop: Robot Component Editor
$3.$ The Arcade: Main Gameplay